Please help me solve the following questions:
a. 1/x+1+2/x-2
b. Write down the sums and products of the roots of the following equations
(a). ax^2-x(a+2)-a=0
(b). X-1/2=3/x+2

Answer by Alan3354(69443)

Please help me solve the following questions:
a. 1/x+1+2/x-2
------
Ambiguous, add parentheses.
--------------
b. Write down the sums and products of the roots of the following equations
(a). ax^2-x(a+2)-a=0
---------------
(b). X-1/2=3/x+2
Ambiguous, add parentheses.